### v1

# Theory of Mind Enhances Collective Intelligence

Abstract
Collective Intelligence plays a central role in a large variety of fields, from economics and evolutionary theory to neural networks and eusocial insects, and it is also core to much of the work on emergence and self-organisation in complex systems theory. However, in human collective intelligence there is still much more to be understood in the relationship between specific psychological processes at the individual level and the emergence of self-organised structures at the social level. Previously psychological factors have played a relatively minor role in the study of collective intelligence as the principles are often quite general and applicable to humans just as readily as insects or other agents without sophisticated psychologies. In this article we emphasise, with examples from other complex adaptive systems, the broad applicability of collective intelligence principles while the mechanisms and time-scales differ significantly between examples. We contend that flexible collective intelligence in human social settings is improved by our use of a specific cognitive tool: our Theory of Mind. We identify several key characteristics of psychologically mediated collective intelligence and show that the development of a Theory of Mind is a crucial factor distinguishing social collective intelligence from general collective intelligence. We then place these capabilities in the context of the next steps in artificial intelligence embedded in a future that includes an effective human-AI hybrid social ecology.
1 Introduction
All intelligence is collective intelligence. [70]
Collectives are capable of achieving things that individuals alone cannot. Notwithstanding the simplicity or complexity of the individuals, their aggregate behaviour can often be understood as a complex processing of information that individuals store, modify, and transfer between each other producing ‘useful’ collective behaviour at the scale of the whole collective. In most instances of Collective Intelligence (CI), where the agents might be ants in an ant colony, bees in a beehive, or neurons in a neural network, the individual is not aware of the drivers of their behaviour or the behaviour of other agents. For example, a single neuron is neither aware of its own internal processes nor that of a neuron it is connected to, nor is it aware of the end goal to which its activity contributes. Despite both this lack of awareness and the lack of a centralised controller, evolutionary and learning processes have produced an intricate, precise, and highly adaptive system that is capable of functional behaviour that would be impossible for any single neuron to achieve. In other instances of CI, such as teams of humans, or businesses interacting in economic markets, the agents themselves may be highly complex and exhibit varying degrees of purposefulness and awareness. Within this context, we draw attention to the role of psychological factors in improving the CI of human social collectives and quantifying the intelligence of social collectives, both natural and artificial.
In order to understand how collectives process information, we first consider the variety of ways in which agents interact. The topology of the network describing agent-to-agent interactions is well known to be important for the proper functioning of social groups [83, 79]. In particular it has been shown that mammalian social groups exhibit patterns of fractal-like topologies [40, 51] that are a result of a cognitive ability to form discrete social connections between conspicifics [49]. These links are often both spatially and temporally transient; people meet for a while, go their separate ways, and come back together later. Despite this transience, individual connections are often the basis of long term social relationships between specific individuals as in pair-bonding and friendships. Consequently an important distinction can be made regarding connections between agents in complex adaptive systems: they can be more fluid-like or more solid-like [101]. For example the links between neurons in the brain are relatively fixed in nature when compared to the brief communicative interactions between ants, either instantaneous interactions between individual ants or via transient pheromone trails that coordinate the behaviour of large numbers of ants. Solé and colleagues [101, 88] identify a distinction between solid brains, in which interactions between agents fixed in place are highly persistent in time (e.g. neural networks, spin glasses) and liquid brains, in which interactions between highly mobile agents are much more short-lived (e.g. ants, immune cells). As Solé et al. note regarding liquid brains [101]: “Here there are no neural-like elements and yet in many ways these systems solve complex problems, exhibit learning and memory, and make decisions in response to environmental conditions.”
All biological agents are composed of sub-units such as organs, cells, and molecular networks [67, 69, 71]. Cells in particular are the simplest living organisms with individual intelligence, or competencies [67, 33], within their native contexts. Here, we briefly focus on the archetypal single-cell intelligence, the neural cells. It is well understood that the central nervous system is a highly developed, adaptive, complex system that exhibits emergent computational characteristics [52], both in biological and artificial neural networks. Naturally the artificial models are simplifications but the extent to which they are simplifications is not so well understood. In a 2021 study, Beniaguev et al. [9] concluded that between five and eight layers of an artificial deep neural network are required to approximate the input–output mapping of a (single) cortical neuron and that the dendritic branches can be understood as spatiotemporal pattern detectors. This demonstrates that a single neural cell can be modelled as an artificial agent with highly complex computational capabilities situated within an adaptive, complex network of other highly complex agents, all signalling to one another. These results can be compared with earlier studies in which neurons were modelled as a Bayesian agent that is trying to infer the state of a hidden variable [25]. In each of these interpretations, a single cell can be seen as an agent with computational competencies situated within the context of a network that is slowly and adaptively changing around it.
We can also compare the competencies of neural cells in networks to the individual competencies of ants in an ant colony. In a recent study [56] it was shown that social structures of some ant colonies are conserved between species that are separated by more than 100 million years of evolution. In the five species studied by Kay et al. [56], they found two social clusters and similarities in the division of labour that are preserved between the species. In a different study, Richardson et al. [91] showed that individuals within an ant colony play an important leadership role and that the behaviour of these individuals significantly improved the collective performance of the ants. Ants are also capable of changing their social structure in the event of pathogenic infestation of their colony. In a 2021 article, Stockmaier and colleagues [102] review the research on social distancing and other social restructuring that occurs with conspecifics in order to reduce the impact of pathogens by changing their social cues, signals, and other behaviours for the collective benefit of the colony. These two very different systems, neural networks and ant colonies, are examples of complex collective intelligences where the individuals (neurons, ants) are complex in their own right, but they signal each other in order to restructure their relationships so as to adapt their collective competencies to external signals. The neural networks are prototypical solid brains and ant colonies are prototypical liquid brains.
Human social interactions can also be viewed as a form of liquid intelligence. Migliano et al. [79] discuss the ‘fluidity’ of social relations in early human societies: “Quantification and mapping of hunter–gatherers’ social networks has revealed details of a fluid and multilevel sociality, where friendship links connect unrelated mobile households into camps of temporary composition”. They describe the key characteristics of early human society, such as egalitarianism, division of labour, cooperative living with unrelated individuals, multi-locality, fluid social structures, and high mobility between campsites, which might be thought of as a liquid brain composed of social interactions that both cluster and disperse in order to store, modify, and transfer information via social networks. The notion that human social interaction might be a form of computation is not new: Mirowski, Axtell and colleagues [82, 60, 81] have suggested that economic markets are a form of computation by which prices can be derived, and Harré recently hypothesised [45] that this could be measured using information theory as had been done earlier for financial markets [47, 42]. As Axtell et al. [60] wrote: “There is a close connection between agent computing in the positive social sciences and distributed computation in computer science, in which individual processors have heterogeneous information that they compute with and then communicate to other processors.”
The emergence of computation in multi-agent systems is a well-studied area of complex adaptive systems [64, 84]. For example neuroscience has used information theory to describe the storage, transfer, and modification of bits of information in biological neural processes [117]. More broadly, Integrated Information Theory (IIT) [107, 77] has been put forward as a measure of the emergence of ‘consciousness’ in generic (non-biological, non-neural) systems. In this case, some forms of IIT explicitly use information theory [8, 78] to measure the amount of non-trivial computation a system is carrying out. More generally, there is a move towards understanding complex adaptive systems in computational terms [89, 74] by empirically measuring the inter-agent flow of information [12].
In this article we use information theory to quantify how much computation in a CI is ‘emergent’ and how much is simply independent information processing by single agents. In general, we wish to capture the notion of the whole (computational process) being greater than the sum of the (independent) parts. We translate this to the simple notion that to the extent to which this inequality holds: Whole - ($\sum$(∑ (Parts)>0)>0) > 0 is the extent to which we will say a system exhibits non-trivial CI, noting that there are multiple possible implementations of this approach [54]. The Parts is how much computation a single agent is carrying out from one time step to the next such that the sum is the total of all agents’ independent computations. The Whole is the totality of computation in the system, it includes all single agent computations, pairwise computations, and higher order interactions between agents. Our measure will not be unique in any of its specifics, but it serves to quantify the CI of a system for comparative analysis. This approach also has much in common with that of Moore et al. [84] in which information theory is used to measure the collective intelligence in biological systems.
Not only is there diversity in the types of systems that can show positive measures of CI, but the ways in which agents manipulate a system’s computations is diverse as well. Take for example Watson and Levin’s discussion of a scientist manipulating the intercellular signalling in order to change their collective outcome [110]:
This framework [of collective cellular intelligence] makes a strong prediction: if intercellular signalling (not genes) is the cognitive medium of a morphogenetic individual, it should be possible to exploit the tools of behavioural and neuro-science and learn to read, interpret and re-write its information content in a way that allows predictive control over its behaviour (in this case, growth and form) without genetic changes.
A counter question is: How can single agents, such as human leaders, have predictive control over a social group? Just as a scientist external to a cell collective can manipulate inter-cellular signalling to control the outcomes of the cell collective, a leader internal to a human collective can manipulate inter-personal behaviours to control the outcomes of the human collective. In both cases, an agent with a goal-directed psychology is acting on inter-agent relationships, i.e. inter-cellular or inter-personal, to control outcomes at the next level higher, i.e. organism-scale or societal-scale.
In this work we will ask an analogous question of human agents: What is there in human psychology that allows us to learn to read, interpret, and re-write our interpersonal information content in a way that allows predictive control over our collective behaviour? We will not be able to explore all of the possible interpretations of this question here, but we posit that our Theory of Mind (ToM) is a suite of cognitive skills that allows individuals to have goal directed control over collective outcomes. Originally ToM was used to describe our ability to infer the unobserved mental states of other people [34] such as desires and beliefs, an ability humans are particularly good at and other animals much less so [86, 59]. But recently it has been shown that ToM is predictive of group performance as well [121, 31], empirically demonstrating the role of ToM in going beyond representations of the internal states of others to using that knowledge in a social setting to improve the collective outcomes for the group. In order to model ToM in a tractable fashion, we will focus on the narrower game theory of mind [123], and the Beliefs, Preferences, and Constraints (BPC) interpretation of game-theoretic decisions put forward by Gintis [35]. In this approach, what agents understand of other agents’ hidden states are the BPC that structure their observable behaviours.
We will consider this question in the framework of agent interactions that extend agent utilities in a simple but novel way. We quantify our results using information theory to show the impact that a correctly deployed ToM has to direct agents’ behaviours in order to increase our CI. The models are simple but they illustrate the central notion that understanding the “beliefs, preferences, and constraints” [36, 37] of others can be used to improve the CI of a complex social system.
In Section 2, we describe the liquid–solid dichotomy of interacting agents, review extant models of ToM, and provide perspective on the interplay between social network structures and ToM. In Section 3, we provide illustrative examples supporting different aspects of our argument, introducing our measure of computation and applying it to a simple empirical example. In Section 4 we review the psychology of social fluidity and the variety of social outcomes that this fluidity makes possible. We also use a simple multi-agent system to describe how a ToM can be used to improve the computational processes, i.e. the CI, of interacting agents. Finally, in Section 5, we discuss the broader implications of this approach.
2 Cognitive morphospaces: The network topology of agents’ interactions
The emergence of cognitive networks marked a pivotal moment in our evolutionary history [120]. Earlier, microorganisms had developed collective structures capable of responding to the physical environment, particularly conditions threatening individual cells [5] and survival became dependent on information exchanges within these groups. Habituation and the ability to minimise the energy response to danger stimuli is considered one of the simplest forms of learning and has been extensively studied in simple collective intelligences such as slime moulds [13, 106]. Similarly bacteria use quorum sensing to coordinate their behaviour based on the density of the population of their community to coordinate responses, leading to a change in gene expression and function regulation e.g. bioluminescence, release of toxins, and biofilm formation [72]. Information processing and problem solving capabilities developed using a variety of interaction types and network topologies long before the appearance of central nervous systems with fixed neuronal structures [5]. But this leads to an interesting question regarding the typologies of agent-to-agent interactions and the intelligence these structures might enable, a question that can be approached by looking at the morphospace of collective intelligence.
A morphospace is a theoretical framework used to simplify and organize the complex shapes and forms of organisms, typically focusing on external anatomical features, into a more manageable space representing their potential variations. For instance, in [3], a three-dimensional morphospace was constructed for organisms with shells, where the diversity of shell shapes is described by three key parameters: a deviation angle, a translation factor, and a growth factor. This reduces a high-dimensional structural space to a lower-dimensional one, where a small number of parameterised properties captures key variations between forms. Morphospaces have found uses in many different fields of research, including the body shapes of fish [20], network topologies [3], and the structural forms of language [96].
These structures, which represent recurring patterns of trait variation, are of great interest to evolutionary biologists because they may indicate shared evolutionary processes and their constraints. They are also of great interest to researchers investigating the underlying structures of collective intelligence [68], as intelligence in its different forms may also be subject to shared evolutionary mechanisms and constraints. With this in mind, work has been done in studying a variety of morphospaces related to collective intelligence and in the next section we review some examples.
2.1 Network topology and the “Solid Brain, Liquid Brain” framework
In the most general of terms, a cognitive network has multiple information processing components that exchange information with each other and interact with their environmental context. Solé et al. [101] identified two key dimensions for categorising different types of cognitive networks: the system’s physical state—either more liquid or more solid in nature, differentiated by how freely individual agents (components) can move in space—and the presence or absence of neurons. The collective dynamics of a large population of agents is influenced by the individuals’ mobility which dictates how they respond to signals both internal and external to the collective. To help conceptualise this diversity, Solé et al. [101] developed a morphospace and taxonomy in order to compare and contrast the physical states of different types of CI. In this way they were able to consider the physical properties that form constraints on the computations achievable by a system. This then poses an interesting question: Is the entire space of possible intelligences being exploited by either synthetic biology or abiotic computation?
Liquid brains exhibit cognitive behaviours without neurons. For example, models by Watson and colleagues [112, 111] illustrated how systems of self-interested agents, driven by a simple mechanism of strengthening beneficial connections, can lead to robust group-level adaptation and problem-solving. This self-organisation, akin to Hebb’s rule in neural networks, enables the system to recall and consequently leverage past configurations that were successful. This also allows the system to generalise from experience and then predict beneficial states it has not encountered before, highlighting how decentralised actions can produce a form of CI that guides the system towards greater global utility. The agents in these models are very simple, but this need not be the case—each agent within a collective may itself have a solid brain as in human social networks where fluid social interactions allow each solid brain to connect and communicate with other solid brains for the benefit of the collective [14]. This is extended to a hybrid model by Kao et al. [55], wherein the modular organisation of mobile animal populations (as an example of a liquid brain) suggests that certain communication pathways exhibit localised and persistent characteristics, akin to those in solid brains. These pathways enhance collective decision-making in complex environments and in turn raise important questions regarding the relative strengths of liquid and solid brains in different contexts. In particular, the conditions under which liquid brains outperform solid brains, especially in terms of adaptability and scalability, remain an open area of investigation. Which computational problems are more effectively solved by liquid brains vis-à-vis solid brains remains to be better understood.
Whole classes of models that describe which biological or artificial structures are capable of some form of computation, either in potentia or in practice, can often be usefully represented using morphospaces [16]. Computational morphospaces [101] have proven effective in studying key properties of complex adaptive systems, for example the statistical mechanics of information processing and structural variations [2, 4]. This sheds light on how energy constraints influence the evolution and adaptability of neural networks across a variety of biological systems. Arsiwalla et al. [2] have examined how liquid brains and solid brains fit within such a framework, by comparing the flexibility and adaptability of different neural architectures [2]. The dimensions of their framework are three different types of complexity that a system may display: autonomic, computational, and social complexity. We suggest that other axes for consideration are a system’s solid–liquid dimension as Solé et al. [101] and Ollé-Vila et al. [85] have done, as well as the system’s degree of ToM (see Section 2.2) and the systems information processing capacity (see Section 3).
We posit that the collective intelligence that emerges from liquid brains (human social networks) is enhanced by our individual capacity for a ToM, where individuals are aware of the goals of others as well as that of the collective, which can then be achieved by adaptation at the local level. Rather than collective intelligence arising as an epiphenomenon or byproduct of agents interacting, our ToM allows agents to causally affect the outcome of the system they are a part of.
2.2 Models of another agent’s internal states
Frith and Frith defined Theory of Mind (ToM) as how we explain other people’s behaviour on the basis of their internal cognitive states, i.e. their knowledge, beliefs, and desires [34]. There is now a vast literature on this topic in psychology, sociology, and more recently artificial intelligence, but for the purposes of this article we restrict ToM to apply to the subset of the beliefs, preferences, and constraints of other agents in the sense of incentivised decisions. This borrows from the BPC model [35] put forward as an approach to understanding the socio-cognitive aspects of human decision-making [37].
One way to interpret the BPC model is that it imposes structural constraints on the process by which decisions are made, and then optimal decisions are discovered within these constraints by parametric variation. Recent work by Peterson et al. [87] compared more than 20 structurally constrained models of individual decision-making using human data. That study was extended to human data during strategic interactions by Harré and El-Tarifi [48] in order to test agents’ constrained representations of other agents. This extends Yoshida et al.’s [123] notion of a Game Theory of Mind to a larger variety of models in which an agent’s strategic reasoning about other agents modulates their behaviour.
In the field of artificial intelligence, Jara-Ettinger [53] proposed the use of Inverse Reinforcement Learning (IRL) as a model of agential ToM, whereby agents modelling other agents’ mental states is equivalent to inferring an unobserved world model the other agent uses in their decision-making, as well as their reward function. Jara-Ettinger discusses some key limitations of IRL, such as the difficulty in recovering an agent’s beliefs and desires even while assuming that all agents are identical in their choice-making. IRL has been implemented in many different algorithmic forms, and their applicability as a basis for ToM was recently reviewed by Ruiz-Serra and Harré [93].
What is missing in the Jara-Ettinger perspective is how different internal models, specifically different drivers of behaviour predicated on the BPC of others, influence how agents interact with one another in a social network. Critically, people use and improve upon these information carrying interactions in their social networks, as discussed next. This requires agents to have more than a representation of the world model used by other agents, it needs to be a social world model of how agents are influenced by their interactions with other agents. Beyond inferring internal states, Shteynberg et al. [97] distinguish between awareness of the self, awareness of the self in relation to others, and an awareness of the collective, which is more than just the sum of the self and others, a type of collective awareness they have called Theory of Collective Mind. In a similar vein, Shum et al. [98] proposed a generative model for understanding multi-agent interactions called Composable Team Hierarchies. This approach used stochastic games in conjunction with multi-agent reinforcement learning in order to infer relationships between agents and predict future behaviours.
In cognitive science, recent progress has been made in identifying levels of ToM ability and placing them in a hierarchical structure (see also Yoshida et al.’s Game Theory of Mind [123] and a recent review by Harré [46]). A cognitive agent with zeroth order ToM attributes no cognitive ability to other agents, whereas first order ToM attributes some cognitive abilities to others, and so on. Here we summarise the orders as described by Lombard and Gärdenfors [76] and use A and B to identify two agents that may or may not have any cognitive ability:
  • Zero order ToM: Both A and B’s behaviour is governed by instinct, reflexes, and conditioning and so direct perception of the agents’ interactions with their environment and each other is all that is needed to understand their behaviour.
  • First order ToM: A and B can be attributed with emotions, attention, desires, intentions, or beliefs, but neither agent attributes these properties to any other agent, including themselves.
  • Second order ToM: A attributes to B internal cognitive states, and that A uses this knowledge to understand B’s behaviour. This is the lowest level at which A attributes hidden (cognitive) variables to B in order to explain the causes of B’s actions, i.e. it abstracts causation away from direct perception of the causes of behaviour.
  • Third order ToM: A attributes to B an understating of A’s internal states. To borrow an example from game theory (stag hunt) and early human society (hunter gatherers), when a hunting party stalks an animal everyone shares a common goal whereby each person A knows that the others are aware of A’s goal, such that this cognitive state will causally inform A’s actions. Lombard and Gärdenfors note that it has not been conclusively demonstrated in nonhuman primates and indicate there are alternative views [39, 63].
  • Fourth order and higher ToM: A has an awareness of at least two mental states, their own and that of B. For example Happé reviewed the evidence and suggested that reflecting on one’s own cognitive state relies on the same neuro-psychological functions as those we use to attribute thoughts to others [41].
Most interactions between agents across all species will be of the zeroth or first order ToM, where neither agent has higher order cognitive states, has no sense of being aware of other agents nor any self-awareness. There is no sharp line that unambiguously distinguishes between stimulus-response, conscious, and self-conscious agents as it is an open area of research and is a multi-dimensional phenomenon [11], and so the conscious status of agents is likely more incremental than the discrete levels of this scale would indicate, but is a useful framing device.
This leads to another core finding related to ToM: the interplay between language and ToM in humans. Chomsky has noted both how well developed this ability is in humans and how communication sits in relation to our mental states [18, p. 10]:
Communication is not a matter of producing some mind-external entity that the hearer picks out of the world, the way a natural scientist could. Rather, communication is a more-or-less affair, in which the speaker produces external events and hearers seek to match them as best they can to their own internal resources. Words and concepts appear to be similar in this regard, even the simplest of them. Communication relies on largely shared cognoscitive powers, and succeeds insofar as similar mental constructs, background, concerns, and presuppositions allow for similar perspectives to be reached. If that is true—and the evidence seems overwhelming—then natural language diverges sharply in these elementary respects from animal communication.
That is to say, effective communication between people requires an encoding-transmission step and a reception-decoding step, both premised on shared cognitive representations in order to be understood, which we contend is made possible via ToM, i.e. a (shared) representation of each other’s hidden cognitive variables.
Recent work has informed Chomsky’s view here, linking language and ToM with our ability to carry out causal reasoning, particularly in social contexts. In Lombard and Gärdenfors they proposed three key hypotheses relating ToM to cognitive structures [76] (quoted, emphasis added):
  • Theory of mind is an integral element of causal cognition;
  • Generally speaking, the more advanced causal cognition is, the more it is dependent on theory of mind; and
  • The evolution of causal cognition depends more and more on mental representations of hidden variables. […] [C]ausal cognition allows us to reason from a network of hidden variables …
Language use is related to this causal reasoning about hidden variables via the representational view of language due to their intersection with ToM [32]. In the representational view of language people use specific grammatical structures to represent complex events and then to reason from them [23]. These structures serve as a cognitive tool, particularly in representing others’ mental states, enabling the expression of false beliefs, lies, or mistakes. It has also been shown that these are strong predictors of children’s false belief understanding, a canonical test of ToM, and targeted syntactical training can improve false belief reasoning [24]. Notably, in developmental learning, language has a stronger influence on ToM abilities than vice versa [80]. From this we identify causal reasoning about hidden cognitive states with causal reasoning about the behaviour of others via a shared representation of our social environment, mediated by specific syntactical structures. We discuss shared representations and their use further in the next section.
2.3 Configurations of our social networks inform individual reasoning
Frith and Frith have previously considered the benefits that accrue to humans via our ToM [34], but what is functionally happening when we use our ToM in social groups? It has been shown that in early hunter-gatherer societies that some emergent phenomena at the social level, e.g. Dunbar’s Number [26, 27], are a consequence of the layered, fractal topology of social networks [28, 40, 51], and that these are in turn the product of very specific, discrete, cognitive constraints at the individual level that shaped the structures of early human societies [49]. At the individual level, a recent review by Momennejad [83] collected the evidence for different social network topologies and how they integrate interpersonal knowledge differently, showing that topology allows social networks to serve a rich variety of collective goals. Momennejad also reviews the neuro-imaging evidence showing that humans neurologically encode these topologies and these encodings are shared across the members of a social group. This was also demonstrated in the work of Lau et al. [65] showing that people are able to integrate information about how agents relate to one another in addition to how they relate to oneself in order to infer social group structures. Lastly, there is evidence for improved collective intelligence when individuals with higher competencies in ToM are present in the group, as shown in the study by Woolley et al. [121]. It was found that, just as there is for an individual person a measure of general cognitive ability, usually denoted $g$, there is an equivalent measure of collective intelligence, denoted $c$, for a group of people. They noted a key explanatory factor of a group’s task performance, as measured by $c$, was the proportion of group members who ranked highly on the Reading the Mind in the Eyes cognitive test introduced by Baron-Cohen and colleagues [6, 7], a test used to measure an individual’s capacity for ToM.
A key takeaway from the work of Woolley and colleagues [121] is that the $c$ factor is not strongly correlated with either the average or maximum intelligence of the individuals in a group. However, it does correlate well with the average social sensitivity of its members, the evenness of the distribution of contributions to group discussions, as well as the proportion of people in the group who rate highly on a ToM test. So there is considerable evidence for the role our ToM plays in group performance, how this shapes interpersonal interactions between people, and, as a consequence, the emergent topological properties of our social groups. This provides support for the argument that ToM is (one of) the individual, bottom-up mechanism(s) through which agents form higher-order social structures with measurable collective intelligence. This may be how we “learn to read, interpret and re-write our interpersonal information content” as we asked following the Watson-Levin quote in Section 1. Next, we illustrate these ideas in a dyadic and then a triadic example of agents interacting with each other and exhibiting non-trivial measure of CI.
3 Topological and cognitive structures in CI: illustrative examples
3.1 A dyadic example of individual learning at short time scales
To illustrate how we will use information theory as a proxy for Woolley et al.’s $c$ intelligence, we use a dataset that was previously studied [44] to measure the information flow in an iterated economic game experiment between monkeys and computers, based on data from an earlier study by Lee et al. [66]. In that study, the experiment had a monkey playing the matching pennies game against a computer algorithm for a reward, the (Nash) optimal reward for the monkey was received if it plays 50:50 across its two choices. A simplified description of the three algorithms used by the computer follows, see Lee et al. [66] for the exact descriptions:
  • Algorithm 0: Play uniformly and independently of the monkey’s choices,
  • Algorithm 1: The computer stores the history of the choices made by the monkey, then to predict what the monkey would do in each trial the computer calculates the conditional probability of the monkey’s choice given the monkey’s choices in the preceding 4 trials, a Null hypothesis was used to test if the monkey played 50:50 and if it was not rejected the computer plays 50:50, if it was rejected the computer plays probability $1-p$ against the monkey’s probability of playing $p$
  • Algorithm 2: Uses the same algorithm as 1) but includes both choices and rewards in the monkey’s history, and then both algorithm 1) and 2) were tested against the Null of 50:50 and if the Null was not rejected the computer plays 50:50, otherwise the best reply was played based on the estimated bias of the monkey.
The usefulness of this example is fourfold. First, it is simple enough that computations can be tested and evaluated in an illustrative way so that the central information theory measures can be applied. Second, it uses game theory as the foundational mechanism for the interactions that generate the information flow between the agents. Third, it is complex enough to illustrate key elements of CI with only two agents. Finally, it illustrates the general principles that can be applied in evolution, psychology, AI collectives, and human-AI hybrid settings. Our definition of CI is based loosely on the information theory form of Integrated Information Theory (IIT) [107], where we are neutral to the interpretation of IIT as a measure of consciousness, and will not be carrying out any optimisation over binary partitions of state variables, so in that sense this is not the same as IIT, although there are some similarities. Our definition for a system with $n$ agents is (see Section 2.2.3 of [78] and for a general introduction see Battencourt [10]):
$\phi(X;\tau)\triangleq I(X_{t};X_{t-\tau})-\sum_{i=1}^{n}I(X^{i}_{t};X^{i}_{t-\tau}).$ (1)
For example we might have $X_{t}=\{X^{1}_{t},X^{2}_{t}\}$ as the joint stochastic variable of agents 1 and 2 such as the monkey and the computer. The function $I(X^{i}_{t};X^{i}_{t-\tau})$ is the time-delayed mutual information (TDMI) with delay $\tau$ for any times series of a stochastic (possibly joint) variable $X_{t}\in\{X_{1},X_{2},\ldots,X_{T}\}$:
$I(X_{t};X_{t-\tau})=\sum_{i=1}^{n}p(X_{t},X_{t-\tau})\log\Big{[}\frac{p(X_{t},X_{t-\tau})}{p(X_{t})p(X_{t-\tau})}\Big{]}.$ (2)
Equation 2 encodes the number of bits that variable $X_{t}$ is able to predict about its own future state based on its ($\tau$-lagged) past states. For example, in a two-agent system, if $X_{t}=\{X^{1}_{t},X^{2}_{t}\}$ then $I(X_{t};X_{t-\tau})$ is the amount of predictive information data $\tau$ steps in the past is encoded in the current state of the entire joint state of the system. On the other hand, if $X_{t}=X^{1}_{t}$ then $I(X_{t};X_{t-\tau})$ encodes how much information the past of one agent encodes about its current behaviour. The difference between the whole system TDMI and the sum of the individual’s TDMI is the extent to which information is being exchanged between the agents, i.e. $\phi(X;\tau)$ in Equation 1 is the excess TDMI.
Table 1 shows the results of these computations. We note the fact that, as the computer algorithm becomes more sophisticated, from algorithm 0 to 1 to 2, $\phi(X;\tau)$ decreases as the monkey plays closer and closer to the Nash strategy, but then as the sophistication increases, the monkey looks further into the past data to extract information. This is most notable in the difference for algorithm 2: at $\tau=3$ (0.0148 bits) is nearly twice that of $\tau=1$ (0.008 bits). This is because, in this specific case, the monkey decouples itself from the computer in order to earn its highest reward: as the computer strategy becomes more sophisticated in measuring the monkeys use of the computer’s choices, the less coupled to the computer and its own past the monkey needs to be.
Time delay: $\tau$ Algo Joint TDMI Monkey TDMI Computer TDMI Excess TDMI: $\phi(X;\tau)$
1 0 0.0999 0 0 0.0999
2 0 0.1197 0.0075 0.0002 0.1120
3 0 0.1237 0.0095 0.0095 0.1047
1 1 0.0693 0.0011 0.0001 0.0681
2 1 0.0748 0.0025 0.0001 0.0722
3 1 0.0766 0.0029 0.0005 0.0732
1 2 0.0105 0.0005 0.0002 0.0080
2 2 0.0155 0.0019 0.0023 0.0113
3 2 0.0216 0.0030 0.0038 0.0148
Table 1: The simplest example of two agents interacting with one another via game theory in which $\phi(X;\tau)$ is non-zero. In general we cannot know the details of the flow of information, only that there is a net positive flow across all agents. All values are in bits and computations carried out using JIDT [75].
3.2 A triadic example of evolutionary learning at long time scales
Before we introduce ToM for social interactions, we consider a second example where evolution has found an agent-to-agent interaction that is similar to the worked example used next in Section 4. Our evolutionary example is a three-agent system: the larval stage of the fly Liriomyza huidobrensis, the pea plant family Fabaceae that fly larvae predate on, and the parasitic wasp Opius dissitus that predates on L. huidobrensis. These species interact in the following way [116]: The larvae of L. huidobrensis infest a pea plant, the pea plant gives off volatiles, called infochemicals, that attract the wasps to the plant, which in turn feed on the larvae, and this larva–wasp conflict indirectly benefits the pea plant.
In this triadic relationship, the pea plant does not directly respond to the threat from the larvae—for example it has not evolved a chemical agent that repels the larvae-laying flies. Instead it signals a third party, the wasp, to bring the wasp into contact with the larvae, and the wasp then eats the larvae. The wasps and the flies are in a dyadic evolutionary competition that could be modelled using two-agent evolutionary game theory, but the pea plant, having facilitated an instance of this conflict, benefits indirectly from the wasps success, in a sense plants can employ other species as a kind of ‘body guard’ [58]. We note that, like the example in Section 4, the plant only signals the wasps to a dyadic interaction when the plant detects the larvae, so that the plant only signals wasps when the plant perceives an intermittent information carrying cue from its environment that wasps are needed.
This type of second order interaction occurs in other ecological examples as well [38] where Trait-Mediated Indirect Interactions (TMIIs) induce hyper-graphs of interactions between species. More complicated interactions have also been observed in which a plant that is being attacked emits infochemicals that lead to unaffected plants emitting volatiles to attract predators [58], to reduce the risk of the unaffected plant being attacked while also aiding the plant being attacked. Kobayashi and Yamamura [58] specifically call this evolutionary development a form of altruism but of course there is no cognitive aspect to this altruism. These examples illustrate that evolution produces forms of sophisticated, inter-species, mixed competitive-altruistic interaction networks but without any need for a ToM, individual awareness, or strategic understanding of the interactions, and so individuals are not knowingly strategic or altruistic as we might interpret a person to be. But these evolved strategies are limited, by definition, to be fixed within the lifetime of a single agent and they can only adapt on evolutionary time-scales. This is in contrast to a ToM which allows us to adapt to multiple strategic and social contexts that may require novel solutions within the lifespan of a single person.
4 Social network modification through ToM: a minimal model
Figure 1: Graphs and hyper-graphs of three agents: (a) A disconnected graph containing two dyadically connected agents and one isolated agent. (b) A connected dyadic graph. (c) A dyadic complete graph. (d) A hyper-graph in which all three agents are connected by a single hyper-link. (e) A combination of a hyper-graph connecting all agents and a disconnected dyadic graph. (f) A connected dyadic graph and a single hyper-link.
In this model we introduce a simple example of interacting agents that form a hypergraph [38]. As in the previous examples, this model is general enough to allow the agents to have any form of biological or artificial psychology, to have zeroth order ToM or fourth order ToM, and it applies to artificial collectives just as readily as it can to biological and ecological collectives.
Game theory provides a formal approach to the analysis of incentivised social interactions in which agents attempt to optimise the outcome (utility) of their joint actions. The expected utilities can be rewritten as polynomials in the agents’ decision variables, usually interpreted as probabilistic weights, $x_{i}\in\mathbf{x}$ for which the co-factors $\mathbf{a}$ are derived from a game payoff matrix (as described in Appendix A). These can be generalised to higher order polynomials representing higher order interactions between the agents: in these expanded utilities the quadratic terms represent the dyadic interactions between agents, the cubic terms represent the three-way interactions, and so on:
[…]
We note that in general $\frac{\partial U_{j}(\mathbf{x};\mathbf{a})}{\partial x_{i}}$ describes the impact that $i$’s choice has on agent $j$’s utility and that we need not assume that there exist either symmetrical or even reciprocal impacts between agents’ utilities and their behavioural choices. In Figure 1, we illustrate how three agents are connected via dyadic relationships and higher order (cubic) hyper-graph interactions. Taking the utility for Agent 1 in Figure 1 as an example, the most general description of the interactions in the dyadic utilities for the top row (a)–(c) are given by:
$(\textrm{a}):\,\,U_{1}(\mathbf{x};\mathbf{a})$ $=a_{1}^{0}+a_{1}^{1}x_{1}$ $\,\,(\textrm{b})\,\textrm{and}\,(\textrm{c}):\,\,U_{1}(\mathbf{x};\mathbf{a})$ $=a_{1}^{0}+\sum_{j}a_{1}^{j}x_{j}+\sum_{j,k}a_{1}^{jk}x_{j}x_{k}$ (4)
In (a), Agent 1 is the only agent that can influence their utility, but for (b) and (c), they are also influenced by other agents, linearly and possibly quadratically, depending on the payoff structure. The co-factors $\mathbf{a}$ of these three utilities are derivable directly from the utility matrices of dyadic agent-to-agent interactions in conventional game theory [50, 43] (see Appendix A). For utilities (d)–(f) in Figure 1:
$(\textrm{d}):\,\,U_{1}(\mathbf{x};\mathbf{a})$ $=\sum_{j,k,l}a_{1}^{jkl}x_{j}x_{k}x_{l}$ $\,\,(\textrm{e}):\,\,U_{1}(\mathbf{x};\mathbf{a})$ $=a_{1}^{0}+a_{1}^{1}x_{1}+\sum_{j,k,l}a_{1}^{jkl}x_{j}x_{k}x_{l}$ (5)
$(\textrm{f}):\,\,U_{1}(\mathbf{x};\mathbf{a})$ $=a_{1}^{0}+a_{1}^{1}x_{1}+\sum_{j,k}a_{1}^{jk}x_{j}x_{k}+\sum_{j,k,l}a_{1}^{jkl}x_{j}x_{k}x_{l}$ (6)
With these descriptions of higher order interactions we can now provide an illustrative example of how a ToM may be used to reconfigure a network to enhance the computational processing of information to improve collective performance. In the example of the parasitic wasp and the fly larvae eating the leaves of a plant, the plant benefits indirectly from the direct interaction between two other agents. In this case the plant encourages the wasp to come into contact with the larvae in a fight for survival that the plant does not directly participate in. In the example that follows next, a similar but theoretical scenario between three agents is studied where two agents are initially interacting but not producing anything with a third agent on the sidelines. Then the third agent provides a signal to the other two to selectively change their behaviour that directly benefits two of the agents and indirectly benefits the third.
4.1 Model scenario
We begin with three agents in proximity to one another situated in a noisy environment in which it is possible for them to collectively do something useful but they are initially in the unfortunate situation in which the behaviour of the three agents produces nothing that is of value. The possible actions ($x_{i}$) of the agents ($A_{i}$) are binary: $x_{i}\in\{-1,1\}$, $i\in\{1,2,3\}$ and so in this example the $x_{i}$ are not probabilities. $A_{1}$ is randomly and uniformly changing its states due to a (useful, information carrying) signal it receives from the environment at time $t$: $s^{t}\in\{-1,1\}$ such that $P(s^{t}=1)=P(s^{t}=-1)=0.5$. $A_{2}$ and $A_{3}$ are initially engaged in the prisoner’s dilemma (PD) game where, as a consequence of selfishly (naïvely) optimising their choice of $x_{i}$, they are in the defect-defect Nash equilibrium (NE), and so their joint action is constant (with zero value). They are capable of a second output when in the cooperate-cooperate configuration, and this output has a positive value. However, being stuck in the PD NE, they are not initially able to produce it.
$A_{2}$ and $A_{3}$ can produce something of value when they cooperate, but it only has value if the external signal $s^{t}=+1$. The output value of $A_{2}$ and $A_{3}$ interacting is initially a sequence of 00s: they are not cooperating with each other, and nothing of value is being produced. We assign each agent a state at time $t$, $x^{t}_{i}\in\{-1,1\}$, such that their dynamic is an ordered sequence of binary states $[x^{1}_{i},x^{2}_{i},\ldots,x_{i}^{T}]$ for $t\in\{1,\ldots,T\}$. The output from $A_{2}$ and $A_{3}$ interacting at time $t$ is a result of having matched their states: $V(o^{t})=1$ if $x^{t}_{2}=x^{t}_{3}=1$ and $V(o^{t})=0$ if $x^{t}_{2}=x^{t}_{3}=-1$, the $x^{t}_{2}\neq x^{t}_{3}$ cases are never achieved as they are not NE and $A_{2}$ and $A_{3}$ will always choose according to the NE of their interactions. The collective behavioural vector is: $\mathbf{x}=\{x_{1},x_{2},x_{3}\}$ or time indexed: $\mathbf{x}^{t}=\{x_{1}^{t},x_{2}^{t},x_{3}^{t}\}$.
Figure 2: The interaction networks for the two scenarios: (a) $A_{1}$ perceives signal $s^{t}$ but cannot relay it to $A_{2}$ and $A_{3}$ and these two agents are not cooperating and so produce 00 value from their output $o^{t}$. (b) $A_{1}$ influences $A_{2}$ and $A_{3}$ to cooperate when $A_{1}$ receives a signal from the environment that cooperating will result in a positive payoff for $A_{2}$ and $A_{3}$; as a consequence, $A_{1}$ receives a portion of the utility from both $A_{2}$ and $A_{3}$.
The signal $s^{t}$ that $A_{1}$ receives indicates whether or not $A_{2}$ and $A_{3}$ should cooperate at time $t$, but initially no information is passing from $A_{1}$ to $\{A_{2},A_{3}\}$ and $A_{1}$ cannot produce anything by itself. So $U_{1}(\mathbf{x};\mathbf{a})=0$, and agents $A_{2}$ and $A_{3}$ only produce output of zero value: $V(o^{t})=0$. The agent utilities in this case are:
$U_{1}(\mathbf{x};\mathbf{a})$ $=0,$ $\,\,U_{2}(\mathbf{x};\mathbf{a})$ $=a_{2}^{0}+a_{2}^{3}x_{3}+a_{2}^{2}x_{2},$ $\,\,U_{3}(\mathbf{x};\mathbf{a})$ $=a_{3}^{0}+a_{3}^{2}x_{2}+a_{3}^{3}x_{3},$ (7)
and $U_{2}=U_{3}=0$ for defect-defect joint strategies. This is equivalent to the configuration in Figure 1(a), and we illustrate this in Figure 2, a. In the second scenario, we assume that $A_{2}$ and $A_{3}$ are still naïvely pursuing their selfish goals, but now $A_{1}$ is more psychologically aware: it knows the beliefs, preferences, and constraints of the other two agents, and it is able to influence their preferences so that they will cooperate with each other when $A_{1}$ receives the signal: $s^{t}=1$. They, in turn, will pay $A_{1}$ some portion of their total payoff.
$A_{3}\rightarrow x_{3}$
$+1\equiv\texttt{C}$ $-1\equiv\texttt{D}$
$A_{2}\rightarrow x_{2}$ $+1\equiv\texttt{Cooperate}$ (R, R) (S, T)
$-1\equiv\texttt{Defect}$ (T, S) (P, P)
(a) Generalised payoff matrix
C D
C $(1,1)$ $(0-c,1+c)$
D $(1+c,0-c)$ $(0,0)$
(b) Prisoner’s Dilemma ($c=-\frac{1}{4}$) and Harmony ($c=\frac{1}{4}$)
Figure 3: Payoff matrices for agents $A_{2}$ and $A_{3}$, which agent $A_{1}$ can strategically influence by setting $c\leftarrow x_{1}$.
To do this numerically, we standardise the symmetric, two-player, two-choice, game theory payoff matrices for $A_{2}$ and $A_{3}$ by setting $R=1$ and $P=0$ (Figure 3, a. and see [43]). We note that the Harmony game (Ha), in which both agents coordinating is the only Nash equilibrium and $T>R>P>S$, can be transformed into the Prisoner’s Dilemma (PD), in which both agents defecting is the only Nash equilibrium and $R>T>S>P$, by introducing a two-state parameter $c\leavevmode\nobreak\ \in\leavevmode\nobreak\ \{-\frac{1}{4},\frac{1}{4}\}\leavevmode\nobreak\ \implies\leavevmode\nobreak\ \{\texttt{PD},\texttt{Ha}\}$ and replacing $T=R+c$ and $S=P-c$ so that the payoff matrices derivable from Figure 3,b. result in the utilities:
$U_{2}(\mathbf{x};\mathbf{a})$ $=R+cx_{2}-(R+c)x_{3},$ $\,\,U_{3}(\mathbf{x};\mathbf{a})$ $=R+cx_{3}-(R+c)x_{2}.$ (8)
As agent $A_{1}$ is aware of the specific BPC model by which $A_{2}$ and $A_{3}$ make their decisions, $A_{1}$ adjusts their behaviour by manipulating their utility co-factors following $s^{t}$, by setting $x^{t}_{1}=\frac{1}{4}s^{t}$ such that $x^{t}_{1}\in\{-\frac{1}{4},\frac{1}{4}\}$. $A_{1}$ then modulates the utility of the other two agents (via an information-carrying signal such as speaking to them) and, in response, receives 10% of the payoffs that result from the interaction between $A_{2}$ and $A_{3}$. The resulting payoffs for each agent are (derived in detail in Appendix B):
$U_{1}(\mathbf{x};\mathbf{a})$ $=\frac{1}{10}\big{(}U_{2}(\mathbf{x};\mathbf{a})+U_{3}(\mathbf{x};\mathbf{a})\big{)},$ $\,\,U_{2}(\mathbf{x};\mathbf{a})$ $=R+x_{1}x_{2}-(R+x_{1})x_{3},$ $\,\,U_{3}(\mathbf{x};\mathbf{a})$ $=R+x_{1}x_{3}-(R+x_{1})x_{2}.$ (9)
4.2 Model interpretation
We make three observations regarding these two scenarios. First, we note that the interactions in the second scenario are a hypergraph in the sense that if any one of the nodes were removed in Figure 2, b. no remaining agent receives any payoff, all agents are necessary contributors to any single agent receiving a utility for their actions [38]. Second, it is straightforward to compute the collective intelligence of both scenarios using Equation 1 and the three stochastic variables $\{x_{1}^{t},x_{2}^{t},x_{3}^{t}\}$ when $s^{t}\in\{-1,1\}$ is uniformly distributed: In the first scenario it is 0 bits and for the second scenario it is 1 bit. This is a simple illustration of how a single agent, knowing the BPC model of two other agents, who are both naïve to all other agents’ BPC models, can improve the collective intelligence of the group by manipulating the network of interactions between agents. This is analogous to the results discussed in Woolley [121] in which group performance, using a measure of collective intelligence, is improved by having agents who have a theory of mind. Third, unlike evolutionary processes, the process in the second scenario has a goal-directed agent who is psychologically informed about the internal states of the other agents and so is able to make plans contingent on different configurations of the interaction network. This implies that $A_{1}$ understands the preferences implied by the co-factors of Equation 8, ($R+c$) and Equation 9, ($R+x_{1}$), influencing the behaviour of the other two agents. These preferences are the hidden (cognitive) variables discussed in Section 2.2, i.e. the Preferences aspect of the BPC framework. In terms of levels of ToM, $A_{1}$ only needs to ascribe hidden states to the other two agents but not the ability for $A_{2}$ or $A_{3}$ to ascribe hidden states to $A_{1}$, for example the NE can be reached by $A_{2}$ and $A_{3}$ by observing the other agent’s behaviour. So the ToM of $A_{1}$ is of second order and $A_{2}$ and $A_{3}$ only need a first order ToM. In this sense a ToM allows $A_{1}$ to reconfigure their social networks to further their goals at much shorter time-scales than the evolutionary time-scale of ecological networks.
4.3 Our perspective in context
We have reviewed some of the recent advances in how people, as complex agents, with a vast space of specialised, context-dependent behaviours, are able to coordinate their activities. In particular, finding ways to recombine our individual competencies to produce, in a short period of time, the appropriate collective competencies is a combinatorically complex task. The central theme of this article is that having a causal (generative) model of another person’s behaviour that reflects the current environmental and social context helps us manage this complexity. This is one way in which our solid brains produce the necessary liquid social structures that quickly build collective solutions with novel collective competencies. In this section we bring these ideas together and summarise our view.
We first draw attention to how, in Section 4, $A_{1}$ has constructed a niche for itself in the context of the preexisting dyadic network between $A_{2}$ and $A_{3}$. In ecological networks a new agent joins a pre-existing network if it can find a niche within which it can fit. This occurs in one of three distinct ways: niche choice, niche conformance, and niche construction [19, 108]. Niche choice occurs when an individual selects environmental conditions that align with its phenotype, while niche conformance involves adjusting its phenotype to suit the environment. Niche construction is the modification of the environment to meet individual needs, which may also impact other species. This suggests an analogy between the formation of ecological networks, where new agents joining may enhance or disrupt the current configuration, and social networks where membership can be explicitly or implicitly gated according to a prospective agent’s ‘fit’ within the group, and even if the group can adjust to accommodate a newcomer, just as the newcomer can adjust in order to be accepted. In human social groups people can be recruited or excluded depending on their contribution to the better functioning of the collective, which in turn corresponds to changes in individual neural activity related to the social network structure [95].
In our analogy with ecological networks, we suggest that any signalling agents in a network need to encode messages over an information-carrying medium that receivers are receptive to and that can then be decoded by a specific receiver. We have argued that humans, as both signallers and receivers, take advantage of our ToM in order to understand what signals will be correctly interpreted by another person and to adapt their signalling to the cognitive state of the receiver. Specifically, our psychology allows us to learn how to read, interpret, and re-write our interpersonal communication over a very short time frame, allowing us rapid, targeted control over our collective behaviour.
In some sense the adage there is nothing new under the sun holds here as evolution and biology have recycled fundamental, pre-existing principles in the service of human sociality. But the adaptive speed of our social networks, the psychological mechanisms involved, the variety of purposes they serve, and the complexity of the communications appear most highly developed in humans. In ecological networks, for example, some agents can act as an encoder-sender of semantic information via infochemicals that signal another agent. A second agent then acts as a receiver-decoder of this signal that in turn changes the behaviour of the receptive agent. The combination of the receiver’s anatomical configuration and behavioural phenotype elicits an appropriate stimulus-response that benefits the signalling agent, and these interactions can form vast, complex hypergraphs of competition and cooperation between agents of many different species. But neither agent needs higher cognitive faculties—their ToM is of the zeroth order. By analogy, humans can target specific individuals or groups of individuals in order to have them adjust their behaviours to best suit the goals of the signaller. The competencies that a ToM affords the sender allows them to know that a receiver is capable of both decoding the signal and acting appropriately in response because they know of the receiver’s receptive and causal states, both cognitive (hidden) and behavioural (overt). That is to say they take advantage of their ToM to understand how, when, what, and to whom they need to signal in order to achieve a beneficial outcome, whereas blind evolution would take much longer to achieve the same results.
The separation of time-scales also plays an important distinction between learning processes that use the same theoretical foundation. The mathematical description of the processes that underpin the evolution of species has been shown to be equivalent to the statistical learning process that underpins Bayesian learning in individual agents [17, 113, 114], for both biological and artificial agents. What Bayesian learning provides at the level of the individual agent is an advantage in the speed of adaptation that Bayesian learning via evolution cannot achieve. Again, this suggests a universality of description that only varies qualitatively in terms of time-scales and the specifics of the mechanisms.
5 Discussion: Theory of Mind in the Context of Social AI
Each step in our cultural development has changed the ways in which we combine individual skills to achieve better, more sophisticated collective outcomes. There is evidence that hunter gatherers participated in the division of labour, complex social networks, multilevel, and fractal-like social structures long before we settled into villages [22, 29, 49, 40]. This was in part due to our ability to construct our own niches [61] and this practice has persisted from hunter-gathers to farming [92] through to civilisation building. As Arroyo-Kalin et al. [1] quote in their opening to the special issue Civilisation and Human Niche Construction:
It is impossible to avoid the conclusion that organisms construct every aspect of their environment themselves. They are not the passive objects of external forces, but the creators and modulators of these forces. The metaphor of adaptation must therefore be replaced by one of construction, a metaphor that has implications for the form of evolutionary theory (Levins and Lewontin 1985: 104).
It has also been argued that through the manipulation of our environmental niche we have brought about the Anthropocene [57, 100, 30, 118].
On the other hand, social niche construction [94] extends this idea to the process whereby agents modify their social context so as to influence their own social evolution. A social niche is the context in which social behaviour occurs, and so social niche construction is where agents actively choose to change their social environments, for example, by choosing who to associate with and how to behave [61]. Ryan et al. [94] describe this in game theoretical terms as the effective game agents are playing after all relevant factors are accounted for, such as the underlying game itself (the payoff matrices in conventional games for example) and any social niche modifiers, amongst others. A social niche modifier is a trait that alters the effective game being played, causing it to differ from the immediate payoff matrices that are usually the complete description of the incentivised interactions, for example population structure, relatedness, punishment etc. In the example in Section 4, $A_{1}$ constructs a social niche for itself by manipulating the structure of the game that $A_{2}$ and $A_{3}$ are playing, for the benefit of $A_{1}$ and incidentally for the benefit of the other agents as well. However, social niche construction theory pertains to evolution more broadly and is not specific to human social networks as the formal description of the model in Section 4 could be an evolved network of plants and animals or a more rapidly changing social network.
In moving towards AI that is situated within a hybrid human-AI ecology, the complexities of effective network construction, communication, and cognitive tools will need to be worked through. Here, we discuss just two of the issues: the difficulties of building psychologically complex AIs and the social environment AI will need to adapt to. The capacity of current AI theories to be sufficient to encompass, in principle at least, human levels of cognition is a rich area of research, from attention mechanisms [109], to reinforcement learning subsuming reasoning [99] and chain of thought reasoning in language models [73] there are arguments being made for emergent phenomena in AI [115] as well as ToM [103] and even artificial general intelligence [15]. What is often missed in these examples is that the individual AI’s ability to digest information and update parameter weights is only a small fraction of what is needed to be effective in a specifically social context. An important tool in our psychological toolbox is our ability to maintain shared hidden variables that are the causal basis of our coordinated joint actions in physical and social environments. Even progress on single agents having effective causal models of the physical environment has been much slower than in other areas of AI [122]. As we have shown in this article, there is very good evidence that causal models, both physical and social, are necessary for people to be able to communicate with each other, and communication is inextricably tied to our ToM. This triad of language, shared causal models of hidden variables, and ToM appears to be a minimum for human social coordination.
Even with this triad established within an AI, the next challenge is where, when, and how an AI should fit into any given collaborative social context. Simple machine intelligence in collaboration with people, such as auto-correct, recommender systems, or GPS navigation, are effective because a human decided there was a need for these tools, they placed them in the appropriate context and then the users adjust their behaviour to any shortcomings the tool might have. But the more autonomous machine intelligence becomes, the more complex the physical and social environment is, and the more trust that needs to be placed in the causal models that drive the behaviour of an AI (i.e. their analogue of a person’s hidden cognitive variables) the more an AI needs to know how to interact with us in a way that resembles how we interact with each other. They will ultimately need to be able to do this via niche construction, niche adaptation, and niche choice, all of which, for people, is a negotiated relationship between each other that needs to be satisficing for those involved. This brings human-AI joint adaptation closer to what Laland et al. have in mind when considering a rethink of evolutionary theory [62]:
We hold that organisms are constructed in development, not simply ‘programmed’ to develop by genes. Living things do not evolve to fit into pre-existing environments, but co-construct and coevolve with their environments, in the process changing the structure of ecosystems.
Consequently, we argue that for any AI to be a beneficially adaptive, autonomous, and socially aware agent, it will also need to reflect the universal principles we see in many evolutionary transitions [105, 104, 90]. This includes multi-level selection in economic [118, 119] and biological systems [21, 71] as well as the hierarchical transitions [21] that appear to be ubiquitous in our biological, social, economic, and technological history.
Appendix A Utility polynomials
In two-player, $m$-action normal-form games, the payoffs are encoded in a single $m\times m$ matrix G, where each cell contains the payoffs for the $i$ and $j$ players for a particular combination of actions.
$x_{j}$
C D
$x_{i}$ C $(g_{i}^{cc},g_{j}^{cc})$ $(g_{i}^{cd},g_{j}^{dc})$
D $(g_{i}^{dc},g_{j}^{cd})$ $(g_{i}^{dd},g_{j}^{dd})$
Table 2: Two-player, two-action normal-form game payoff matrix G
The payoff matrix G can be decomposed into two standard matrices $\texttt{G}_{i},\texttt{G}_{j}$ containing each player’s respective payoffs. Utility functions can be obtained in polynomial form from the elements $g_{i}^{l}\in\texttt{G}_{i}$, where each term intuitively represents the contribution of each agent’s action (first-order terms) and their interaction (second-order terms) [43]. For $m=2$ actions, we have:
$U_{i}(\mathbf{x};\mathbf{a}_{i})=a_{i}^{0}+a_{i}^{i}x_{i}+a_{i}^{j}x_{j}+a_{i}^{ij}x_{i}x_{j}$ (A1)
$\mathbf{a}_{i}=\begin{bmatrix}a_{i}^{0}\\ a_{i}^{i}\\ a_{i}^{j}\\ a_{i}^{ij}\end{bmatrix}=\frac{1}{4}\begin{bmatrix}g_{i}^{cc}+g_{i}^{cd}+g_{i}^{dc}+g_{i}^{dd}\\ (g_{i}^{dc}+g_{i}^{dd})-(g_{i}^{cc}+g_{i}^{cd})\\ (g_{i}^{cd}+g_{i}^{dd})-(g_{i}^{cc}+g_{i}^{dc})\\ (g_{i}^{cc}+g_{i}^{dd})-(g_{i}^{cd}+g_{i}^{dc})\end{bmatrix}$ (A2)
A.1 General form for $n$ players and $m=2$ actions
For any number $n$ of players where each player’s action $x_{i}\in\{-1,1\}$, the utility function is
$U_{i}(\mathbf{x};\mathbf{a}_{i})=\sum_{S\subseteq\{1,2,\ldots,n\}}a_{i}^{S}\prod_{j\in S}x_{j}$ (A3)
where $S$ is any subset of the set $\{1,2,\ldots,n\}$, and $a_{i}^{S}$ is the co-factor corresponding to the subset $S$. For example, $S=\{1,2,3\}$ means the product $\prod_{j\in S}x_{j}=x_{1}x_{2}x_{3}$, and $a_{i}^{S}=a_{i}^{1,2,3}$ represents the co-factor for this interaction. To determine the co-factors $a_{i}^{S}$, we need to consider the payoff matrix $\texttt{G}_{i}$, which contains $2^{n}$ entries corresponding to each combination of $n$ players’ actions. Denote the entry in $\texttt{G}_{i}$ for the action combination $(x_{1},x_{2},\ldots,x_{n})$ as $g_{i}^{x_{1}x_{2}\ldots x_{n}}$. The co-factors $a_{i}^{S}$ are:
$a_{i}^{S}=\frac{1}{2^{n}}\sum_{(x_{1},x_{2},\ldots,x_{n})\in\{-1,1\}^{n}}g_{i}^{x_{1}x_{2}\ldots x_{n}}\prod_{j\in S}x_{j}$ (A4)
This generalization for the utility function for $n$-player, two-action games can be interpreted through Fourier analysis on the Boolean cube. This connection arises because we represent the utility function as a sum of basis functions over the Boolean domain, where the co-factors $a_{i}^{S}$ are Fourier co-factors and $\prod_{j\in S}x_{j}$ are parity functions. Thus, we express the utility function $U_{i}$ as a Fourier transform of the payoff matrix $\texttt{G}_{i}$. Each co-factor $a_{i}^{S}$ captures the influence of a subset $S$ of players on the overall utility, analogous to how Fourier co-factors capture the influence of different frequency components in a signal. This generalization leverages the principles of Fourier analysis on the Boolean cube to decompose the complex payoff interactions into simpler, orthogonal components, making it easier to analyze and interpret the contributions of different action combinations in the game.
Appendix B Derivation of utility polynomials in example
From the provided payoff matrix,
C D
C $(1,1)$ $(0-c,1+c)$
D $(1+c,0-c)$ $(0,0)$
we can obtain the polynomial co-factors (see Table 2 and Equation A2 for details)
$\mathbf{a}_{i}=\begin{bmatrix}a_{i}^{0}\\ a_{i}^{i}\\ a_{i}^{j}\\ a_{i}^{ij}\end{bmatrix}$ $=\frac{1}{4}\begin{bmatrix}g_{i}^{cc}+g_{i}^{cd}+g_{i}^{dc}+g_{i}^{dd}\\ (g_{i}^{dc}+g_{i}^{dd})-(g_{i}^{cc}+g_{i}^{cd})\\ (g_{i}^{cd}+g_{i}^{dd})-(g_{i}^{cc}+g_{i}^{dc})\\ (g_{i}^{cc}+g_{i}^{dd})-(g_{i}^{cd}+g_{i}^{dc})\end{bmatrix}$ (B1)
$=\frac{1}{4}\begin{bmatrix}1+(0-c)+(1+c)+0\\ ((1+c)+0)-(1+(0-c))\\ ((0-c)+0)-(1+(1+c))\\ (1+0)-((0-c)+(1+c))\end{bmatrix}=\frac{1}{4}\begin{bmatrix}1+1\\ 1+c-1+c\\ -c-1-1-c\\ 1+c-1-c\end{bmatrix}$ (B2)
$=\frac{1}{4}\begin{bmatrix}2\\ 2c\\ -2(1+c)\\ 0\end{bmatrix}=\frac{1}{2}\begin{bmatrix}1\\ c\\ -(1+c)\\ 0\end{bmatrix}=\begin{bmatrix}a_{i}^{0}\\ a_{i}^{i}\\ a_{i}^{j}\\ a_{i}^{ij}\end{bmatrix}$ (B3)
and substitute into the (second-degree) utility polynomial (with zero quadratic term in this particular case)
$U_{i}(\mathbf{x};\mathbf{a})=a_{i}^{0}+a_{i}^{i}x_{i}+a_{i}^{j}x_{j}+a_{i}^{ij}x_{i}x_{j}=\frac{1}{2}\Big{(}1+cx_{i}-(1+c)x_{j}\Big{)}$ (B4)
we can choose to re-scale the utility values by a factor of 2, which results in the following utility polynomials for $A_{2}$ and $A_{3}$, respectively
$U_{2}(\mathbf{x};\mathbf{a})$ $=1+cx_{2}-(1+c)x_{3},$ $\,\,U_{3}(\mathbf{x};\mathbf{a})$ $=1+cx_{3}-(1+c)x_{2}.$ (B5)